OpenStudy (anonymous):

simplify 4^3 square root of 3x + 5^3 square root of 10x

OpenStudy (anonymous):

\[4^3 \sqrt{3x} + 5^3\sqrt{10x}\] This?

OpenStudy (anonymous):

\[4^3\sqrt{3x} +5^3\sqrt{10x}\]

OpenStudy (anonymous):

lol yep

OpenStudy (anonymous):

Correctomundo =P

OpenStudy (anonymous):

\[ (64 \sqrt{3}+125 \sqrt{10}) \sqrt{x} \] Simplified formalities want you to have "constant*variables"

OpenStudy (anonymous):

Really though it's just bureaucracy formality for this problem, and practice with factoring out exponents. \[\sqrt{x}=x^{1/2}\]

OpenStudy (anonymous):

How sweet of them to want me to have constant variables.